# Rethinking the Evaluation of Dialogue Systems: Effects of User Feedback on Crowdworkers and LLMs

### v1

@@ Abstract.
In ad-hoc retrieval, evaluation relies heavily on user actions, including implicit feedback. In a conversational setting such signals are usually unavailable due to the nature of the interactions, and, instead, the evaluation often relies on crowdsourced evaluation labels. The role of user feedback in annotators’ assessment of turns in a conversational perception has been little studied. We focus on how the evaluation of task oriented dialogue systems, is affected by considering user feedback, explicit or implicit, as provided through the follow-up utterance of a turn being evaluated. We explore and compare two methodologies for assessing TDSs: one includes the user’s follow-up utterance and one without. We use both crowdworkers and the large language models as annotators to assess system responses across four aspects: relevance, usefulness, interestingness, and explanation quality. Our findings indicate that there is a distinct difference in ratings assigned by both annotator groups in the two setups, indicating user feedback does influence system evaluation. Workers are more susceptible to user feedback on usefulness and interestingness compared to LLMs on interestingness and relevance. User feedback leads to a more personalized assessment of usefulness by workers, aligning closely with the user’s explicit feedback. Additionally, in cases of ambiguous or complex user requests, user feedback improves agreement among crowdworkers. These findings emphasize the significance of user feedback in refining system evaluations and suggest the potential for automated feedback integration in future research. We publicly release the annotated data to foster research in this area We publicly release the annotated data 111 https://github.com/Clemenciah/LLMCrowdDialogueEval/tree/main/Data.
@@ 1. Introduction
Evaluation of systems has been an integral part of the information retrieval (IR) research agenda for decades (e.g., Cleverdon et al., 1966). Traditionally, IR evaluation has relied highly on user actions, including implicit feedback such as click-through rates. However, in a conversational setting such signals are not usually available due to the nature of the interactions. As a result, the evaluation of dialogue systems increasingly relies on human evaluation which has led to a growing interest in user-centric evaluation methods (Zhang and Balog, 2020). However, asking for explicit user feedback from a user can be intrusive and may negatively impact user experience (Zhang et al., 2018). Therefore, in recent years, the assessment of conversational systems has relied on crowdsourced evaluation, leveraging the collective wisdom of human annotators.
Turn-level assessments. When gathering evaluation feedback on individual turns in a conversational interaction, various design methods have been considered in the past. These include deciding on the type of judgment scale, as well as formulating annotation guidelines and methods for presenting dialogues under assessment at the turn level (Roitero et al., 2018; Siro et al., 2024). Recent strategies for presenting turn-level utterances to annotators involve two main approaches: one displays both the user’s initial request and the system’s response for evaluation (Sun et al., 2021; Siro et al., 2023), and the other shows only the system’s response (Mehri and Eskenazi, 2020). The choice between these approaches often depends on the specific evaluation metric in question. The first method, similar to the query–document pair setting in ad-hoc retrieval systems, operates under the premise that the user’s initial request offers sufficient context for annotators to make well-informed evaluations.
@@
We use a subset of the recommendation dialogues (ReDial(Li et al., 2018) dataset to address the following research questions:
  1. RQ1
    How does a user’s implicit feedback from the follow-up utterance influence the evaluation labels collected from both human annotators and LLMs?
  2. RQ2
    When is implicit user feedback significant in the evaluation of TDS?
  3. RQ3
    What are the annotators’ perceptions in terms of the sources of information they rely on to make their assessments, and what are the potential biases might that have on their performance?
Findings. Our findings indicate that both the crowdworkers and the LLM exhibit sensitivity to user cues from follow-up utterances. There is a significant difference in the mean ratings from both annotators except for relevance when follow-up utterance is included indicating user feedback does influence system evaluation. Workers are more susceptible to user feedback in usefulness and interestingness compared to LLMs in interestingness and relevance. Specifically, there is a clear distinction in relevance and usefulness ratings of crowdworkers Setup 2 ratings unlike in Setup 1 where these aspects are often conflated. This indicates that crowdworkers not only evaluate response usefulness based on topical relevance but also alignment with user needs and preferences expressed in follow-up utterance. This suggests that follow-up utterances enable a more personalized assessment of usefulness, aligning closely with the user’s explicit feedback. In Setup 2, we observe an increase in annotator agreement. This is particularly evident in scenarios characterized by uncertainty in user requests. Complex user requests with multiple criteria or preferences posed challenges during evaluation, but follow-up utterances helped to clarify the user intent. Similarly, generic user requests, initially broad and challenging to address, became more focused on follow-up utterances, allowing human/LLM-based annotators to tailor their assessment effectively.
@@ 2. Related Work
Recent studies emphasize evaluating TDSs through a user experience lens (Deriu et al., 2020). Traditionally, TDSs were assessed primarily for task completion. While task completion remains a fundamental criterion, there is a growing recognition that solely measuring task success may not provide a comprehensive understanding of system performance (Siro et al., 2023). Consequently, there is a shift towards incorporating user experience metrics into the evaluation of TDS (Siro et al., 2022). Here, we provide a brief overview of the studies on TDS evaluation from the perspective of user feedback, evaluation bias, and LLMs.
@@ 2.2. Bias in crowdsourcing evaluation labels
The use of crowdsourcing for evaluating TDS in IR research, while offering scalability and diversity, brings inherent biases. Research, such as (Eickhoff, 2018; Santhanam et al., 2020; Azzopardi, 2021), highlights cognitive biases and cognitive load as significant factors influencing crowdworker judgments, which can skew the evaluations. For example, workers’ preconceived notions or mental fatigue might lead to inconsistent results. Further, Hube et al. (2019) and Han et al. (2020) emphasize the impact of workers’ relevance strategies and personal biases on their assessment of IR systems. To mitigate these biases, strategies such as task design adjustments and worker training are suggested (Gadiraju et al., 2017). For instance, presenting tasks neutrally and providing clear, unbiased instructions can help reduce bias, as discussed in (Hube et al., 2019). Additionally, the choice and implementation of judgment scales, as explored in (Barbera et al., 2020), play a crucial role in assessor bias. Different from previous studies, in this work we focus on assessing how the sources of information relied on by workers to make their judgment bias their evaluation labels.
@@ 3. The Annotation Task
Our objective is to understand the influence of user feedback from the user’s follow-up utterance on the evaluation of TDSs. We conduct our study as an annotation effort with crowdworkers from amazon mechanical turk (AMT(Amazon Mechanical Turk, 2023). Additionally, due to the increased use of LLMs as annotators (He et al., 2023; Ding et al., 2023) we seek to understand how LLMs are affected by feedback from the user’s next utterance. User feedback in this case can be either implicit or explicit. Explicit feedback refers to straightforward, direct responses from users, like specific comments on a certain dialogue aspect (e.g., “I don’t like this movie”). Implicit feedback is more subtle, encompassing aspects like tone or contextual hints within the user’s follow-up utterance (e.g., “Thanks for the suggestion. How about some action movies?”). We gather annotations for four fine-grained dialogue qualities: relevance, usefulness, interestingness and explanation quality, across two experimental conditions.
@@ 3.1. Dialogue qualities
Explanation quality. Understanding how well a TDS communicates its reasoning is important for user trust and comprehension. Explainability in IR systems has a witnessed a notable surge in recent times (Guo et al., 2023; Balog and Radlinski, 2020; Zhang et al., 2014; Zhang and Chen, 2020). Following Guo et al. (2023) we instruct workers to assess explanation quality, by evaluating the clarity and informativeness of the system’s justifications or explanations accompanying its responses. This aspect provides insights into the system’s transparency and user-friendly communication.
Interestingness. Beyond system functionality, the interestingness (Siro et al., 2022) of a system response adds a subjective layer to the evaluation. Workers are asked to evaluate whether the system’s responses are engaging, captivating, or exhibit qualities that make the interaction more enjoyable for the user. This encapsulates the language used to make recommendation by the system. This aspect contributes to a holistic assessment of user experience.
@@ 3.3. Annotation scale
Following the approach outlined by (Roitero et al., 2018), our study adopts the S100 scale for evaluation purposes. This scale is employed through a sliding window mechanism, allowing annotators to provide detailed feedback on the dialogue systems. The sliding scale’s interactive nature enables a more precise and flexible assessment compared to traditional binary or categorical scales. To enhance its usability and ensure intuitive responses, the default value on this slider is set to 0. This design choice is based on the rationale that a neutral starting point encourages annotators to consciously adjust the slider based on their judgment of the dialogue turn’s effectiveness, rather than being biased by any preset values. To ensure consistency and accuracy in evaluations, we provide the annotators with several examples demonstrating how to effectively use the S100 scale. We adopt the same scale for annotation with LLM.
@@ 3.4. Preliminary experiments
  1. (1)
    Single worker, two conditions (Exp 1): Workers evaluated a dialogue turn under two conditions within a single human intelligence task (HIT). The only difference was the presence or absence of the user’s follow-up utterance, which served as user feedback.
  2. (2)
    Random assignment of conditions (Exp 2): Workers were randomly assigned to one of the two conditions to incorporate diverse perspectives and reduce potential biases, gathering a range of rationales behind annotator evaluations.
@@ 3.5. Experimental conditions
Following Exp 2 we designed two distinct experimental conditions to evaluate the effect of user feedback on the evaluation of TDSs with human annotators, as well as LLMs.
@@ 1. Setup 1
    Following the conventional method of annotation, this condition provides only the initial user query and the system’s response to the annotators and LLMs, omitting the user’s follow-up utterance. This setup focuses on evaluating the TDS based on a single interaction, reflecting the traditional approach in dialogue annotation.
@@ 3.7. LLM as annotator
We used ChatGPT (gpt-3.5-turbo API  [^2] Temperature = 1, Top p = 1) a subseries of GPT models (Brown et al., 2020) for the annotation task. Typically, crowdworkers are provided with annotation examples to improve the quality of the labels they assign, similarly, we provide the same to ChatGPT, thus using it in a few-shot setting. To ensure consistency, we replicate the experimental conditions used for human annotators across the four aspects. In our experiments, we employ two distinct experimental conditions, varying the prompts  [^3] An online appendix with the prompts and additional analysis is available at https://github.com/Clemenciah/LLMCrowdDialogueEval/tree/main/Appendix and data presentation to align with these setups. For each aspect, we provided ChatGPT with the corresponding human annotation instructions. This approach allows us to comprehensively assess the performance of ChatGPT in generating evaluation labels in a manner that mirrors human annotation practices.
@@ 4. Crowdsourced Judgments
Before addressing our research questions (in Section 57), we examine the judgments collected through crowdsourcing.
Internal agreement. To assess the quality of labels collected from the crowdworkers, we compute pairwise Cohen’s Kappa and report the results in Table 1. The Kappa scores indicate varying levels of agreement in different evaluation setups and aspects:
@@
Figure 2. A comparison of individual worker score distributions for Setup 1 (left column) and Setup 2 (right column).
@@
The interestingness aspect is highly subjective, therefore making it more prone to individual worker bias as observed with the moderate agreement between workers in Table 1. In Setup 1 the scores are skewed towards the left, indicating most workers found the system responses less interesting (see Fig. 2(f)) with a median score of 37 compared to 50 for Setup 2.
@@ 5. Effect of user feedback
In this section, we answer RQ1: How does user feedback from the follow-up utterance influence the evaluation labels collected from both crowdworkers and LLMs?
Distributions. For the crowdsourced labels, each turn is annotated by three workers; their ratings are averaged to get the overall score per turn. Fig. 3 (a)–(d) show the density distributions of the scores: Relevance and usefulness are skewed towards the right (Fig. 3(a) and 3(b)) for both setups, showing that more turns are found to be more relevant and useful by the crowdworkers. For usefulness, the peak for Setup 2 is towards the center compared to Setup 1, indicating a decrease in the number of turns that are highly useful as workers have access to the user’s follow-up utterance, adding more context during the assessment. Cases where the system makes a relevant recommendation, are rated highly useful in Setup 1, but the user’s feedback in Setup 2 changes the worker’s rating to lower values in certain cases. E.g., in cases where the user has already watched the movie or even though the movie satisfies their requirements (e.g., genre and actor), they do not like other aspects. Interestingness has a more central peak with a wide range showing there was a lot of variability in the assessment (Fig. 3(c)). More turns are assessed as interesting in Setup 2 compared to Setup 1, with fewer turns being scored as highly interesting in both setups (80–100). Similar observations pertain to explanation quality (Fig. 3(d))
@@
Humans vs. LLM. The different peaks between the two setups across the four aspects indicate a significant divergence in how crowdworkers and LLMs perceive and rate the aspects in different setups. Setup 1 is characterized by high peaks in high-range scores, while Setup 2 exhibits peaks in the moderate range except for relevance, which has both moderate and high peaks. This contrast suggests that the inclusion of the follow-up utterance has a notable impact on both the crowdworkers and LLM assessments.
@@
Agreement with expert ratings. Here, we examine how well human and LLM labels align with expert ratings. We collect expert ratings on the user satisfaction aspect to investigate the correlation of the fine-grained aspects to overall user satisfaction following Siro et al. (2023). Using the same setup, we collect the expert ratings from two experts. Since our initial annotation was on a S100 scale we transform the labels to S3 scale (1–3) for all aspects (Han et al., 2019b) and then calculate the Spearman’s $r$ between each aspect and the expert rating. We report our results in Table 2.
Table 2. Spearman’s $r$ correlation coefficient between the aspects and expert user satisfaction ratings for both the crowdworkers and LLM. Italicized values indicate non-significant values ( $p<0.05$).
@@
Humans vs. LLM. Correlations with expert ratings highlight that relevance and usefulness assessments generally have stronger alignment with expert user satisfaction ratings across setups and annotator types. However, interestingness shows weaker correlations, indicating potential challenges in assessing this aspect consistently and objectively (Siro et al., 2022). In general, we note that humans perform well in assessing user experience measures such as usefulness and interestingness while LLMs perform well in assessing utility measures such as relevance and explanation quality.

### v2

@@ Abstract.
In ad-hoc retrieval, evaluation relies heavily on user actions, including implicit feedback. In a conversational setting such signals are usually unavailable due to the nature of the interactions, and, instead, the evaluation often relies on crowdsourced evaluation labels. The role of user feedback in annotators’ assessment of turns in a conversational perception has been little studied. We focus on how the evaluation of task oriented dialogue systems, is affected by considering user feedback, explicit or implicit, as provided through the follow-up utterance of a turn being evaluated. We explore and compare two methodologies for assessing TDSs: one includes the user’s follow-up utterance and one without. We use both crowdworkers and large language models as annotators to assess system responses across four aspects: relevance, usefulness, interestingness, and explanation quality. Our findings indicate that there is a distinct difference in ratings assigned by both annotator groups in the two setups, indicating that user feedback does influence system evaluation. Workers are more susceptible to user feedback on usefulness and interestingness compared to LLMs on interestingness and relevance. User feedback leads to a more personalized assessment of usefulness by workers, aligning closely with the user’s explicit feedback. Additionally, in cases of ambiguous or complex user requests, user feedback improves agreement among crowdworkers. These findings emphasize the significance of user feedback in refining system evaluations and suggest the potential for automated feedback integration in future research. We publicly release the annotated data.111 https://github.com/Clemenciah/LLMCrowdDialogueEval/tree/main/Data
@@ 1. Introduction
Evaluation of systems has been an integral part of the information retrieval (IR) research agenda for decades (e.g., Cleverdon et al., 1966). Traditionally, IR evaluation has relied highly on user actions, including implicit feedback such as click-through rates. However, in a conversational setting such signals are not usually available due to the nature of the interactions. As a result, the evaluation of dialogue systems increasingly relies on human evaluation, leading to a growing interest in user-centric evaluation methods (Zhang and Balog, 2020). However, asking for explicit user feedback from a user can be intrusive and may negatively impact user experience (Zhang et al., 2018). Therefore, in recent years, the assessment of conversational systems has relied on crowdsourced evaluation, leveraging the collective wisdom of human annotators.
Turn-level assessments. When gathering evaluation feedback on individual turns in a conversational interaction, various design methods have been considered in the past. These include deciding on the type of judgment scale, as well as formulating annotation guidelines and methods for presenting dialogues under assessment at the turn level (Roitero et al., 2018; Siro et al., 2024). Recent strategies for presenting turn-level utterances to annotators involve two main approaches: one displays both the user’s initial request and the system’s response for evaluation (Sun et al., 2021; Siro et al., 2023), and the other shows only the system’s response (Mehri and Eskénazi, 2020). The choice between these approaches often depends on the specific evaluation metric in question. The first method, similar to the query–document pair setting in ad-hoc retrieval systems, operates under the premise that the user’s initial request offers sufficient context for annotators to make well-informed evaluations.
@@
We use a subset of the recommendation dialogues (ReDial(Li et al., 2018) dataset to address the following research questions: (RQ1) How does a user’s implicit feedback from the follow-up utterance influence the evaluation labels collected from both human annotators and LLMs? (RQ2) When is implicit user feedback significant in the evaluation of TDSs? (RQ3) What are the annotators’ perceptions in terms of the sources of information they rely on to make their assessments, and what are the potential biases might that have on their performance?
Findings. Our findings indicate that both the crowdworkers and the LLM exhibit sensitivity to user cues from follow-up utterances. There is a significant difference in the mean ratings from both annotators except for relevance when follow-up utterance is included, indicating user feedback does influence system evaluation. Workers are more susceptible to user feedback in usefulness and interestingness, compared to LLMs in interestingness and relevance. Specifically, there is a clear distinction in relevance and usefulness ratings of crowdworkers Setup 2 ratings unlike in Setup 1 where these aspects are often conflated. This indicates that crowdworkers not only evaluate response usefulness based on topical relevance but also align with user needs and preferences expressed in follow-up utterances. This suggests that follow-up utterances enable a more personalized assessment of usefulness, aligning closely with the user’s explicit feedback. In Setup 2, we observe an increase in annotator agreement. This is particularly evident in scenarios characterized by uncertainty in user requests. Complex user requests with multiple criteria or preferences posed challenges during evaluation, but follow-up utterances helped to clarify the user intent. Similarly, generic user requests, initially broad and challenging to address, became more focused on follow-up utterances, allowing human-/LLM-based annotators to tailor their assessment effectively.
@@ 2. Related Work
Recent studies emphasize evaluating TDSs through a user experience lens (Deriu et al., 2020). Traditionally, TDSs were assessed primarily for task completion. While task completion remains a fundamental criterion, there is a growing recognition that solely measuring task success may not provide a comprehensive understanding of system performance (Siro et al., 2023). Consequently, there is a shift towards incorporating user experience metrics into the evaluation of TDSs (Siro et al., 2022). Here, we provide a brief overview of the studies on TDS evaluation from the perspective of user feedback, evaluation bias, and LLMs.
@@ 2.2. Bias in crowdsourcing evaluation labels
The use of crowdsourcing for evaluating TDSs in IR research, while offering scalability and diversity, brings inherent biases. Research, such as (Eickhoff, 2018; Santhanam et al., 2020; Azzopardi, 2021), highlights cognitive biases and load as significant factors influencing crowdworker judgments, which can skew the evaluations. For example, workers’ preconceived notions or mental fatigue might lead to inconsistent results. Further, Hube et al. (2019) and Han et al. (2020) emphasize the impact of workers’ relevance strategies and personal biases on their assessment of IR systems. To mitigate these biases, strategies such as task design adjustments and worker training are suggested (Gadiraju et al., 2017). For instance, presenting tasks neutrally and providing clear, unbiased instructions can help reduce bias, as discussed in (Hube et al., 2019). Additionally, the choice and implementation of judgment scales, as explored in (Barbera et al., 2020), play a crucial role in assessor bias. Different from previous studies, in this work we focus on assessing how the sources of information relied on by workers to make their judgment bias their evaluation labels.
@@ 3. The Annotation Task
Our objective is to understand the influence of user feedback from the user’s follow-up utterance on the evaluation of TDSs. We conduct our study as an annotation effort with crowdworkers from amazon mechanical turk (AMT(Amazon Mechanical Turk, 2023). Additionally, due to the increased use of LLMs as annotators (He et al., 2023; Ding et al., 2023) we seek to understand how LLMs are affected by feedback from the user’s next utterance. User feedback in this case can be either implicit or explicit. Explicit feedback refers to straightforward, direct responses from users, like specific comments on a certain dialogue aspect (e.g., “I don’t like this movie”). Implicit feedback is more subtle, encompassing aspects like tone or contextual hints within the user’s follow-up utterance (e.g., “Thanks for the suggestion. How about some action movies?”). We gather annotations for four fine-grained dialogue qualities: relevance, usefulness, interestingness, and explanation quality, across two experimental conditions.
@@ 3.1. Dialogue qualities
Explanation quality. Understanding how well a TDS communicates its reasoning is important for user trust and comprehension. Explainability in IR systems has witnessed a notable surge in recent times (Guo et al., 2023; Balog and Radlinski, 2020; Zhang et al., 2014; Zhang and Chen, 2020). Following Guo et al. (2023) we instruct workers to assess explanation quality, by evaluating the clarity and informativeness of the system’s justifications or explanations accompanying its responses. This aspect provides insights into the system’s transparency and user-friendly communication.
Interestingness. Beyond system functionality, the interestingness (Siro et al., 2022) of a system response adds a subjective layer to the evaluation. Workers are asked to evaluate whether the system’s responses are engaging, captivating, or exhibit qualities that make the interaction more enjoyable for the user. This encapsulates the language used to make recommendations by the system. This aspect contributes to a holistic assessment of user experience.
@@ 3.3. Annotation scale
Following the approach outlined in (Roitero et al., 2018), our study adopts the S100 scale for evaluation purposes. This scale is employed through a sliding window mechanism, allowing annotators to provide detailed feedback on the dialogue systems. The sliding scale’s interactive nature enables a more precise and flexible assessment compared to traditional binary or categorical scales. To enhance its usability and ensure intuitive responses, the default value on this slider is set to 0. This design choice is based on the rationale that a neutral starting point encourages annotators to consciously adjust the slider based on their judgment of the dialogue turn’s effectiveness, rather than being biased by any preset values. To ensure consistency and accuracy in evaluations, we provide the annotators with several examples demonstrating how to effectively use the S100 scale. We adopt the same scale for annotation with LLM.
@@ 3.4. Preliminary experiments
  1. Exp 1
    Single worker, two conditions: Workers evaluated a dialogue turn under two conditions within a single human intelligence task (HIT). The only difference was the presence or absence of the user’s follow-up utterance, which served as user feedback.
  2. Exp 2
    Random assignment of conditions: Workers were randomly assigned to one of the two conditions to incorporate diverse perspectives and reduce potential biases, gathering a range of rationales behind annotator evaluations.
@@ 3.5. Experimental conditions
Following Exp 2, we designed two distinct experimental conditions to evaluate the effect of user feedback on the evaluation of TDSs with human annotators, as well as LLMs.
@@ 1. Setup 1
    Following the conventional annotation method, this condition provides only the initial user query and the system’s response to the annotators and LLMs, omitting the user’s follow-up utterance. This setup focuses on evaluating the TDS based on a single interaction, reflecting the traditional approach in dialogue annotation.
@@ 3.7. LLM as annotator
We used ChatGPT (gpt-3.5-turbo API [^2] Temperature = 1, Top p = 1) a subseries of GPT models (Brown et al., 2020) for the annotation task. Typically, crowdworkers are provided with annotation examples to improve the quality of the labels they assign, similarly, we provide the same to ChatGPT, thus using it in the few-shot setting. To ensure consistency, we replicate the experimental conditions used for human annotators across the four aspects. In our experiments, we employ two distinct experimental conditions, varying the prompts and data presentation to align with these setups. [^3] The prompts are available in our Online Appendix in Sec. A. For each aspect, we provided ChatGPT with the corresponding human annotation instructions. This approach allows us to comprehensively assess the performance of ChatGPT in generating evaluation labels in a manner that mirrors human annotation practices.
@@ 4. Crowdsourced Judgments
Before addressing our research questions (in Sec. 57), we examine the judgments collected through crowdsourcing.
Internal agreement. To assess the quality of labels collected from the crowdworkers, we compute pairwise Cohen’s Kappa and report the results in Tab. 1. The Kappa scores indicate varying levels of agreement in different evaluation setups and aspects:
@@
Figure 2. A comparison of individual worker scores distributions for Setup 1 (left column) and Setup 2 (right column).
@@
The interestingness aspect is highly subjective, therefore making it more prone to individual worker bias as observed with the moderate agreement between workers in Tab. 1. In Setup 1 the scores are skewed towards the left, indicating most workers found the system responses less interesting (see Fig. 2(f)) with a median score of 37 compared to 50 for Setup 2.
@@ 5. Effect of user feedback
In this section, we answer (RQ1): How does user feedback from the follow-up utterance influence the evaluation labels collected from both crowdworkers and LLMs?
Distributions. For the crowdsourced labels, each turn is annotated by three workers; their ratings are averaged to get the overall score per turn. Fig. 3 (a)–(d) show the density distributions of the scores: Relevance and usefulness are skewed towards the right (Fig. 3(a) and 3(b)) for both setups, showing that more turns are found to be more relevant and useful by the crowdworkers. For usefulness, the peak for Setup 2 is towards the center compared to Setup 1, indicating a decrease in the number of turns that are highly useful as workers have access to the user’s follow-up utterance, adding more context during the assessment. Cases where the system makes a relevant recommendation, are rated highly useful in Setup 1, but the user’s feedback in Setup 2 changes the worker’s rating to lower values in certain cases. E.g., in cases where the user has already watched the movie or even though the movie satisfies their requirements (e.g., genre and actor), they do not like other aspects. Interestingness has a more central peak with a wide range showing there was a lot of variability in the assessment (Fig. 3(c)). More turns are assessed as interesting in Setup 2 compared to Setup 1, with fewer turns being scored as highly interesting in both setups (80–100). Similar observations pertain to explanation quality (Fig. 3(d)).
@@
Humans vs. LLM. The different peaks between the two setups across the four aspects indicate a significant divergence in how crowdworkers and LLMs perceive and rate the aspects in different setups. Setup 1 is characterized by high peaks in high-range scores, compared to Setup 2 which exhibits peaks in the moderate range except for relevance, which has both moderate and high peaks. This contrast suggests that user feedback from the follow-up utterance has a notable impact on both the crowdworkers and LLM assessments.
@@
Agreement with expert ratings. Here, we examine how well human and LLM labels align with expert ratings. We collect expert ratings on the user satisfaction aspect to investigate the correlation of the fine-grained aspects to overall user satisfaction following Siro et al. (2023). Using the same setup, we collect the expert ratings from two experts. Since our initial annotation was on an S100 scale we transformed the labels to S3 scale (1–3) for all aspects (Han et al., 2019b) and then calculated the Spearman’s $r$ between each aspect and the expert rating. We report our results in Tab. 2.
Table 2. Spearman’s $r$ correlation coefficient between the aspects and expert user satisfaction ratings for both the crowdworkers and LLM. * indicate non-significant values ($p<0.05$).
@@
Humans vs. LLM. Correlations with expert ratings highlight that relevance and usefulness assessments generally have a stronger alignment with expert user satisfaction ratings across setups and annotator types. However, interestingness shows weaker correlations, indicating potential challenges in assessing this aspect consistently and objectively (Siro et al., 2022). In general, we note that humans perform well in assessing user experience measures such as usefulness and interestingness while LLMs perform well in assessing utility measures such as relevance and explanation quality.
Further analysis of the correlation of the aspects in the two setups and task duration is available in our Online Appendix.